OpenStudy (anonymous):

Just need to make sure my work is right. Will give medal to first reply. Multiple choice.

OpenStudy (anonymous):

The same amount of trash is dumped into a landfill every day. The function below shows the total number of tons of trash n in the landfill after x days. n = 2000x + 1000 What does the number 2000 represent? 1. Amount of trash originally in the landfill 2. Amount of trash added to the landfill every day 3. Rate at which the total trash increases every day 4. Rate at which the total trash is recycled every day

OpenStudy (anonymous):

I think its the first one.

OpenStudy (lyrae):

I'd say it's #2

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

@agent0smith can you help me with one more?

OpenStudy (agent0smith):

1000 is the amount originally in the landfill, since when you put in x=0 you'd get n=1000 2000 is the amount added each day

OpenStudy (anonymous):

okay, can you help me with a different problem?

OpenStudy (agent0smith):

k

OpenStudy (anonymous):

Which of the following best defines 3 to the power of 2 over 3? Square root of 9 Cube root of 9 Cube root of 3 Square root of 3

OpenStudy (agent0smith):

\[\Large 3^{\frac{ 2}{ 3 }} = (3^2)^{1/3} = \sqrt[3]{3^2}\]

OpenStudy (anonymous):

okay, thanks